OpenStudy (please_help):

How is doing operations (adding, subtracting, multiplying, and dividing) with rational expressions similar to or different from doing operations with fractions? Can understanding how to work with one kind of problem help understand how to work another type? When might you use this skill in real life?

OpenStudy (espex):

Basically any time you have a value that is dependent on the ratio of two other values. A very simple example from my life experience is finding voltage across different branches of a circuit. V=I/R, so \[V _{t}=\frac{I _{1}}{R _{1}}+\frac{I _{2}}{R _{2}}+\frac{I _{3}}{R _{3}}\]